OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation. x^2 + x - 12 = 0

OpenStudy (anonymous):

y = x^2 + x - 12 = 0 y = (x - 3)(x + 4)

OpenStudy (anonymous):

Can also get if you assume (x+a)(x+b) then -12 = a b +1 = a + b Simple choices for -12 = -(12)(1) = -(6)(2) = -(4)(3) and a=-3 b=4 does the trick.

OpenStudy (anonymous):

alternatively, use the quadratic equation solution formula

OpenStudy (anonymous):

which gives you x = 3, b=-4 so (x-3)(x+4) is factoring